A particle with a charge of -1.24 x 10 C is moving with instantaneous velocity = (1.49 x 101)+(-3.85 x 101. What is the force exerted on this particle by a magnetic field (a) = (1.40 7) + (-2.30 7) and (b) B = (-1.25T) + (2.007)j + (-3.50 T)k?


1
Expert's answer
2021-05-27T18:38:13-0400

At first, we have:


"q=1.24\times10^{-8}\,C; \,\vec{v}=\begin{bmatrix}
1.49\times 10^1, & -3.85\times 10^1, & 0,
\end{bmatrix}

\tfrac{m}{s}"


"(a)\,\vec{B} = \begin{bmatrix}
1.40, & -2.30, & 0,
\end{bmatrix} T"

"(b)\,\vec{B} = \begin{bmatrix}
-1.25, & 2.00, & -3.50,
\end{bmatrix}T"


"\vec{F} = q\cdot \vec{v}\times \vec{B}"


Using the definition of the Lorentz force will allow us to calculate the force exerted by each magnetic field, thus we find that:


(a) "\vec{F_{B}}=(1.24\times10^{-8}) \begin{pmatrix}
\,\widehat{i} & \widehat{j} & \widehat{k}\,\\ 
14.9 & -38.5 & 0\\ 
1.40 &-2.30 & 0
\end{pmatrix}\,N"


"\vec{F_{B}}=(1.24\times10^{-8}\,N)\begin{pmatrix}
14.9 & -38.5\\ 
1.40 &-2.30
\end{pmatrix}\widehat{k}"


"\vec{F_{B}}=(1.24*10^{-8}\,N)((-14.9*2.30)-(-1.40*38.5))
\widehat{k}"


"(a) \vec{F_{B}}=(2.43\times10^{-7}\,N)\,\widehat{k}"


Then, for b:


(b) "\vec{F_{B}}=(1.24\times10^{-8}) \begin{pmatrix}
\widehat{i} & \widehat{j} & \widehat{k}\\ 
14.9 & -38.5 & 0\\ 
-1.25 &-2.00 & -3.50
\end{pmatrix}\,N" =


"(1.24)(10^{-8})\,N)\left [ \begin{pmatrix}
14.9 & -38.5 \\ 
-1.25 & -2.00
\end{pmatrix}\widehat{k}\,+\begin{pmatrix}
\widehat{i} & \widehat{j} \\ 
14.9 & -38.5
\end{pmatrix}(-3.50) 

\right ]"


="(1.24*10^{-8})[(-29.8-48.125)\widehat{k}\,+(-38.5\widehat{i}-14.9\widehat{j})(-3.50)]"


= "(134.75\widehat{i}+52.15\widehat{j}-77.925\widehat{k})(1.24*10^{-8})\,N"


(b) "\vec{F_{B}}=(16.71\widehat{i}+6.47\widehat{j}-9.66\widehat{k}\,) *10^{-7}\,N"


In conclusion, the forces exerted on the charged particle due to the magnetic field are:

  • "(a)\,\vec{F_{B}}=(2.43*10^{-7}\,N)\,\widehat{k}"
  • "(b)\,\vec{F_{B}}=(16.71\widehat{i}+6.47\widehat{j}-9.66\widehat{k}\,) *10^{-7}\,N"


Reference:

  • Young, H. D., & Freedman, R. A. (2014). Sears and Zemansky's University Physics (with Modern Physics Technology Update). Edinburgh: Pearson.